CTS involve defining problems.
A learner adds two fractions as follow :2\3 +6\5 =7/9.
define the problem which you have to work on with this learner.
1
Expert's answer
2020-07-07T20:13:40-0400

We know "\frac{2}{3} + \frac{6}{5}" = "\frac{2}{3}\times \frac{5}{5} + \frac{6}{5} \times \frac{3}{3} = \frac{10}{15} + \frac{18}{15}=\frac{10+18}{15}=\frac{28}{15}" "\neq \frac{7}{9}" .

Hence, the given statement is false.